[Brahms-l] Shift report 20050522 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 22 2005 - 16:05:40 EDT
Shifter leader was kh
Shifters were 


Had beam during the whole day except for 1.5hr between stores.  
Achieved runs 14840 through 14850 starting at 9:30.
There were no obvious detector problems.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 22 16:05:56 2005

This archive was generated by hypermail 2.1.8 : Sun May 22 2005 - 16:06:08 EDT